OpenStudy (anonymous):

simplify 5/6 - 1/2 (1/3 + 3/8)

OpenStudy (anonymous):

help

OpenStudy (anonymous):

Work inside the bracket first. (1/3 + 3/8) This works out to 11/24 (the LCM is 24 i.e. the denominators -- 3 X 8). Next multiply 11/24 by 1/2 = 11/48. Then subtracting 11/48 from 5/6. we get 5/6 = 40/48 - 11/48 = 29/48 which is ur ansa!